OpenStudy (anonymous):

How do I break this up into partial functions? (s^4+81)/ s(s^2+9)^2 Not sure what i should be splitting it up to

OpenStudy (anonymous):

I get someone like A/s + B/(s^2+9) + c/(s^2+9)^2

OpenStudy (amistre64):

a/s b/(s^2+9) c/(s^2+9)^2 where a b and c are the appropriate poly substitutions

OpenStudy (amistre64):

s^2 requres the ax+b linear top for completeness

OpenStudy (amistre64):

same with the other quad, needs a linear atop it

OpenStudy (anonymous):

Ugh So I was kinda right? Not sure how to finish that. Do I solve for ABC, then split up the equation with B in it?

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

\[\frac{a}{s}+\frac{bx+c}{(x^2+9)}+\frac{dx+e}{(x^2+9)^2}\] lordamercy fortunately b = c = 0

OpenStudy (anonymous):

\[a(x^2+9)^2+(bx+c)x(x^2+9)+(dx+c)x=x^4+81\]

OpenStudy (anonymous):

replace x by 0, get \[81a=81\] \[a=1\] so that part was easy

OpenStudy (anonymous):

So I got a=1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now i think you have to expand and equate like coefficients

OpenStudy (anonymous):

So i just need to find d and e right?or is there more to it?

OpenStudy (anonymous):

assuming we grant that b and c are both zero, yes but i am not sure we get to assume that

OpenStudy (anonymous):

whole thing is annoying, but if you just need the answer it is here http://www.wolframalpha.com/input/?i=%28s^4%2B81%29%2F+%28s%28s^2%2B9%29^2%29

OpenStudy (anonymous):

that seems obsurd..

OpenStudy (amistre64):

hmm, did you do a fancy substitution with s=x for the others ? :)